OpenStudy (anonymous):

Mark is in training for a marathon race. He has been training by running 10 miles each day. He will be increasing his running by this weekly formula: . Using the formula, what distance, in miles, will Mark be running in 10 weeks? Round to the nearest mile.

OpenStudy (mayankdevnani):

y = 10(1.1)x, where x = 0 y = 10(1.1)^0 = 10(1) = 10

OpenStudy (anonymous):

Mark is in training for a marathon race. He has been training by running 10 miles each day. He will be increasing his running by this weekly formula: . Compute the distance Mark will run if he continues using the formula, for 52 weeks (x = 52). Does this distance seem reasonable?